### v2

# On Geometric Structures for Policy Parameterization in Continuous Control

Abstract
Standard stochastic policies for continuous control often rely on ad-hoc boundary-enforcing transformations (e.g., tanh) which can distort the underlying optimization landscape and introduce gradient pathologies. While alternative parameterizations on the unit manifold (e.g., directional distributions) are theoretically appealing, their computational complexity (often requiring special functions or rejection sampling) has limited their practical use. We propose a novel, computationally efficient action generation paradigm that preserves the structural benefits of operating on a unit manifold. Our method decomposes the action into a deterministic directional vector and a learnable concentration scalar, enabling efficient interpolation between the target direction and uniform noise on the unit manifold. This design can reduce policy head parameters by nearly 50% (from $2d$ to $d+1$) and maintains a simple $O(d)$ sampling complexity, avoiding costly sampling procedures. Empirically, our method matches or exceeds state-of-the-art methods on standard continuous control benchmarks, with significant improvements (e.g., +37.6% and +112%) on high-dimensional locomotion tasks. Ablation studies confirm that both the unit-norm normalization and the adaptive concentration mechanism are essential to the method’s success. These findings suggest that robust, efficient control can be achieved by explicitly respecting the structure of bounded action spaces, rather than relying on complex, unbounded distributions. Code is available in supplementary materials.
1 Introduction
Continuous control (OpenAI et al., 2019) remains one of the most challenging problems in reinforcement learning (RL) (Silver et al., 2014), with applications ranging from robotics to autonomous driving (Seo et al., 2025). At the heart of this challenge lies a fundamental design choice: how should agents generate continuous actions? For over a decade, Gaussian policies have served as the default answer, powering algorithms from Deep Deterministic Policy Gradient (DDPG) (Lillicrap et al., 2015) to Soft Actor-Critic (SAC) (Haarnoja et al., 2018) and achieving remarkable success across diverse domains. Their popularity stems from mathematical convenience, including closed-form entropy, straightforward reparameterization, and well-understood optimization properties.
Yet this convenience masks a fundamental mismatch. Physical systems operate within bounded action spaces, while Gaussian distributions have infinite support Nikishin et al. (2021). The standard solution applies squashing functions like $\tanh$ to map samples into bounded regions (Theile et al., 2024), but this transformation distorts the distribution’s geometry, creates gradient flow issues near boundaries, and breaks the natural symmetry of the action space (Fujimoto et al., 2018). As policies become more deterministic during training, actions cluster near boundaries where $\tanh$’s gradient vanishes. We observe this phenomenon in $\approx$~40% of SAC training steps on HalfCheetah (Figure A.2, Appendix B). Such instabilities are often misattributed to insufficient exploration, while in fact they reflect a deeper geometric mismatch between Gaussian policies and bounded action spaces.
Recent work has begun questioning this Gaussian orthodoxy Davidson et al. (2022). von Mises-Fisher (vMF) distributions offer a mathematically principled alternative by operating directly on the unit sphere, naturally respecting bounded constraints (Michel et al., 2024). However, vMF’s theoretical elegance comes at a steep computational cost (You et al., 2025): sampling requires rejection methods with $O(dk)$ complexity where $k$ is the expected number of rejections, and density computation involves modified Bessel functions prone to numerical overflow (Mazoure et al., 2019). Other alternatives like normalizing flows or mixture distributions add expressiveness but compound the computational burden (Obando-Ceron et al., 2024). This creates a dilemma: accept Gaussian’s geometric limitations or pay the price of computational complexity.
We take a different path. Rather than seeking increasingly sophisticated distributions, we ask whether the distribution paradigm itself is necessary. Actions in physical systems naturally decompose into direction and magnitude. For instance, a robot arm moves toward a target direction with some force, and a car steers at an angle with some acceleration. This geometric intuition suggests that effective action generation might not require explicit probability modeling at all.
This insight leads to Geometric Action Control (GAC), which generates actions through direct geometric operations on the unit sphere. GAC represents policies through two components: a direction network that outputs unit vectors indicating preferred action orientations, and a concentration network that controls exploration by interpolating between deterministic directions and uniform spherical noise. This decomposition transforms the complex problem of sampling from sophisticated distributions into simple linear interpolation, reducing computational complexity from $O(dk)$ to $O(d)$ while maintaining the geometric consistency that bounded action spaces demand.
Our key contributions are:
  • We introduce GAC, a distribution-free action generation paradigm that replaces probabilistic sampling with direct geometric operations on the unit sphere, challenging the necessity of distributional modeling in continuous control.
  • We develop a compact and efficient policy architecture requiring only d+1d{+}1 parameters instead of $2d$ for Gaussian policies, achieving comparable or better performance with reduced complexity.
  • We provide theoretical analysis showing that spherical mixing achieves vMF-like concentration without Bessel functions, and ablations demonstrating that spherical geometry and adaptive concentration are critical to GAC’s success.
  • We provide comprehensive empirical evaluation across six MuJoCo benchmarks, achieving the best performance on 4 out of 6 tasks, especially in high-dimensional control, validating that geometric consistency outweighs distributional sophistication.
Beyond immediate performance gains, GAC represents a conceptual shift in how we approach policy design. By demonstrating that geometric structure can replace distributional complexity, we open new avenues for developing efficient, interpretable, and theoretically grounded control algorithms. We believe our results champion a broader “Geometric Simplicity Principle”: that for many robotics and control tasks, explicitly modeling the geometric structure of the action space is a more effective and efficient path forward than pursuing ever more sophisticated probability distributions. The remainder of this paper is organized as follows: Section 2 reviews related work, Section 3 presents the GAC methodology, Section 4 provides empirical evaluation, and Section 5 concludes with discussions of broader implications.
2 Related Work
2.1 Gaussian Policies and Their Limitations
The dominance of Gaussian policies in continuous control traces back to the natural policy gradient literature, where Gaussian distributions provided tractable gradient estimates and convergence guarantees. Modern deep RL algorithms like SAC, Proximal Policy Optimization (PPO) (Schulman et al., 2017), and Trust Region Policy Optimization (TRPO) (Schulman et al., 2015) inherit this choice, implementing Gaussian policies through neural networks that output mean and variance parameters. While this approach has driven impressive empirical success, practitioners have long recognized its limitations in bounded action spaces. The standard $\tanh$ squashing solution, popularized by SAC, maps Gaussian samples to bounded intervals but introduces well-documented issues: gradient vanishing near boundaries, asymmetric action distributions, and the fundamental contradiction of using infinite-support distributions for finite spaces (Bendada et al., 2025). Despite these known problems, the field has accepted them as necessary trade-offs for computational convenience.
2.2 Beyond Gaussian: Alternative Distributions
[…]
The most principled alternative emerged from directional statistics (Sinii et al., 2024). vMF distributions, operating directly on the unit sphere, elegantly address boundary constraints through their geometric formulation. Recent work (Scott et al., 2021) demonstrated vMF policies could match or exceed Gaussian performance while providing theoretical advantages. However, vMF’s practical adoption faces significant hurdles (Banerjee et al., 2005): sampling requires rejection methods with acceptance rates as low as 0.1 for high concentrations, likelihood computation involves modified Bessel functions $I_{v}(\kappa)$ prone to numerical overflow for large $\kappa$, and the concentration parameter lacks intuitive interpretation for practitioners (Zaghloul & Johnson, 2025). These challenges have confined vMF policies largely to theoretical investigations rather than practical deployment.
2.3 Geometric Perspectives in RL
Parallel to distributional innovations, a geometric understanding of RL has emerged (Hu et al., 2022). Hyperbolic RL leverages non-Euclidean geometry for hierarchical representations, while work on Riemannian policy optimization extends natural gradients to curved manifolds (Nickel & Kiela, 2017; Wang et al., 2024; Müller & Montúfar, 2024). These approaches highlight how geometric structure can inform algorithm design, yet they typically add complexity rather than simplifying existing methods. Our work inverts this trend, using geometric insights to simplify rather than sophisticate.
The connection between action spaces and geometric structures has appeared implicitly in several contexts (Wang et al., 2025). Quaternion representations for rotational actions in robotics naturally embrace spherical geometry, while circular distributions model periodic actions in locomotion tasks (Zhou et al., 2019). However, these remain specialized solutions rather than general principles. GAC unifies previously scattered geometric insights into a single, practical framework for continuous control—demonstrating that the central question is not which distribution to adopt, but whether explicit distributions are necessary at all.
2.4 Simplification as Innovation
The evolution from TRPO (Schulman et al., 2015) to PPO (Schulman et al., 2017) exemplifies a crucial pattern in deep RL: dramatic simplification often yields superior practical performance. Where TRPO required complex conjugate gradient procedures and line searches, PPO achieved comparable or better results through simple clipped objectives. Similarly, Twin Delayed Deep Deterministic policy gradient (TD3) (Fujimoto et al., 2018) simplified DDPG’s actor-critic architecture while improving stability and performance by 30% on average. These trends suggest that algorithmic complexity in RL often reflects a lack of structural clarity, rather than a theoretical necessity.
GAC follows this simplification philosophy. Rather than adding sophistication to handle Gaussian limitations or implementing complex vMF sampling, we identify the minimal geometric structure necessary for effective control. This approach aligns with recent trends toward interpretable and efficient RL, where understanding why methods work matters as much as empirical performance.
2.5 The Missing Perspective: Action Generation Without Distributions
Most approaches to continuous control operate within the distributional paradigm: policies are defined as probability densities over actions, requiring likelihood evaluation, entropy regularization, and absolute continuity (Engstrom et al., 2020). This perspective, inherited from supervised learning and classical statistics, may be unnecessarily restrictive for control, where actions are ultimately deterministic functions of states and randomness.
GAC replaces this distributional machinery with a geometric operation. Instead of modeling a density over $\mathbb{R}^{d}$, we generate actions via a direction sampled on the unit sphere and a scalar magnitude. This reframes control not as modeling a distribution, but as directly generating structured actions. By replacing distributional complexity with geometric operations, GAC eliminates density evaluations, reparameterization tricks, and explicit entropy calculations while avoiding gradient saturation from $\tanh$ squashing. Recent theoretical work (Tiwari et al., 2025) shows that RL trajectories tend to concentrate on low-dimensional manifolds, using complex mathematical analysis to uncover this emergent structure. GAC inverts the perspective: rather than discovering manifolds, we build on them. By constraining actions to the unit sphere, we achieve structure by design, not by accident. From emergent complexity to designed simplicity—GAC exemplifies a principle in RL: structure need not emerge; it can be constructed.
3 Problem Formulation and Methodology
3.1 Problem Formulation
We consider the standard continuous control setting where an agent interacts with an environment through bounded continuous actions. The action space $\mathcal{A}\subseteq[-1,1]^{d}$ represents normalized physical constraints, where $d$ is the action dimension. The agent observes states $s\in\mathcal{S}$ and selects actions according to a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$.
The maximum entropy RL framework augments the standard objective with an entropy term to encourage exploration: $J(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(r_{t}+\alpha\mathcal{H}(\pi(\cdot|s_{t}))\right)\right],$ where $\gamma\in[0,1)$ is the discount factor, $r_{t}=r(s_{t},\mathbf{a}_{t})$ denotes the reward at time step $t$, $\alpha>0$ is the temperature parameter controlling exploration-exploitation trade-off, $\mathcal{H}$ denotes entropy, and $\tau=(s_{0},\mathbf{a}_{0},s_{1},\mathbf{a}_{1},\dots)$ represents trajectories sampled from the policy-environment interaction.
Figure 1: Architecture of GAC. State $s$ is processed by a shared backbone, which branches into a direction head producing a unit vector $\bm{\mu}$, and a concentration head predicting $\kappa$. The final action is generated via spherical mixing, replacing traditional distributional sampling with direct geometric interpolation.
The Geometric Mismatch. Most continuous control methods model policies as Gaussian distributions $\mathcal{N}(\mu(s),\Sigma(s))$ with unbounded support, requiring $\tanh$ squashing to map samples into $1,1]^{d}$. While this approach has proven highly successful, as evidenced by SAC remaining a leading method, it creates a fundamental mismatch: unbounded distributions must be compressed into bounded action spaces. The $\tanh$ transformation achieves this compression but induces gradient saturation when $|\tilde{\mathbf{a}}_{i}|$ is large, with our analysis showing substantial pre-squashed samples fall in low-gradient regions (Appendix B). SAC addresses this through strong entropy regularization ($\alpha\approx 0.2$), maintaining exploration despite reduced gradients. GAC takes a different path: rather than mitigating the mismatch through entropy-driven exploration, we eliminate it by operating directly on the unit sphere. This geometric approach ensures consistent gradient flow and enables adaptive exploration through learned concentration, offering a structurally simpler alternative that aligns policy support with environmental constraints by design.
3.2 Geometric Action Generation
Core Insight. GAC replaces traditional action sampling with a geometric pipeline consisting of direction mapping, concentration control, and spherical mixing, as illustrated in Figure 1. Rather than modeling probability distributions over actions, we directly generate actions via geometric operations on the unit sphere, a natural choice that aligns with high-dimensional concentration phenomena where directions carry the primary semantic information. This shift eliminates the need for density-based computations such as log-probabilities or entropy, while preserving exploration through structured spherical noise, modulated by a learnable concentration parameter $\kappa$.
Direction Mapping. A neural network $f_{\mu}:\mathcal{S}\rightarrow\mathbb{R}^{d}$ produces raw directional vectors, which are normalized to the unit sphere:
[…]
where $||\cdot||_{2}$ denotes the L2 norm. This normalization ensures $\bm{\mu}(s)\in\mathbb{S}^{d-1}$, the unit sphere in $d$ dimensions.
Concentration Control. A separate network $f_{\kappa}:\mathcal{S}\rightarrow\mathbb{R}$ predicts concentration scores, which modulate the trade-off between deterministic direction and stochastic noise. These scores are transformed via a sigmoid function $w(\kappa)=\sigma(\kappa)\in(0,1)$ to produce the mixing weight used in (2), enabling smooth interpolation and stable, adaptive exploration throughout training.
Spherical Mixing. Actions are generated by interpolating between the deterministic direction and uniform spherical noise:
$\mathbf{a}=r\cdot\text{normalize}\left(w(\kappa)\cdot\bm{\mu}+(1-w(\kappa))\cdot\bm{\xi}\right),$ (2)
where $\bm{\xi}\sim\text{Uniform}(\mathbb{S}^{d-1})$ is sampled as normalized Gaussian noise to provide isotropic exploration on the unit sphere, and $r=2.5$ is the action radius (task-dependent). Even with substantial noise contribution (e.g., 30% when $\kappa\approx 1$), actions remain coherent: spherical normalization preserves directionality while preventing magnitude corruption, ensuring stable control throughout training.
Intrinsic Exploration. In contrast to conventional approaches where exploration is externally injected (e.g., Gaussian noise or entropy bonuses), GAC inherently encodes stochasticity within the action generation process. The random direction $\bm{\xi}$ is not an auxiliary perturbation but an integral part of the policy’s structure, making exploration an intrinsic geometric property. The learnable concentration $\kappa$ acts as an endogenous control signal, adaptively modulating the exploration-exploitation trade-off without external regularization. This eliminates the need for separate entropy bonuses or temperature scheduling (e.g., $\alpha$ tuning in SAC).
Gradient Flow. All GAC operations, including normalization, sigmoid mixing, and linear interpolation, are differentiable and compatible with standard backpropagation, ensuring stable policy optimization.
3.3 Theoretical Justification
The spherical mixing operation creates an implicit distribution with geometrically intuitive concentration control. While a closed-form density is intractable, we rigorously establish how the mixing weight controls distribution concentration.
Theorem 1 (Expected Direction Control).
For GAC’s spherical mixing operation, the expected unnormalized sample vector lies precisely along the mean direction, scaled by the mixing weight:
$\mathbb{E}_{\bm{\xi}}[\mathbf{v}]=w(\kappa)\bm{\mu},$ (3)
where $\mathbf{v}=w(\kappa)\bm{\mu}+(1-w(\kappa))\bm{\xi}$ is the unnormalized mixture, $\bm{\mu}\in\mathbb{S}^{d-1}$ is the mean direction, $\bm{\xi}\sim\text{Uniform}(\mathbb{S}^{d-1})$ is uniform spherical noise, and $w(\kappa)=\sigma(\kappa)$ is the mixing weight with sigmoid function $\sigma(x)=1/(1+e^{-x})$.
This result is exact and requires no approximation: $w(\kappa)$ directly controls the expected alignment with $\bm{\mu}$, and as $\kappa$ grows large ($w(\kappa)\to 1$), variance shrinks and samples concentrate tightly around $\bm{\mu}$. This provides vMF-like concentration behavior without Bessel function computations. See Appendix A.1 for the proof.
3.4 Integration with SAC
GAC naturally integrates into the SAC framework by replacing the standard Gaussian policy with our geometric action generator. The key distinction lies in exploration: GAC eliminates explicit probability computations and entropy regularization, achieving exploration instead through geometric mixing controlled by $\kappa$.
3.4.1 Exploration Control Mechanism
GAC introduces a learned exploration controller $\kappa(s)$ that adaptively modulates the balance between deterministic actions and stochastic exploration. Unlike SAC’s temperature parameter $\alpha$ which requires manual tuning or scheduling, $\kappa$ learns directly from the value landscape.
In the maximum entropy framework, the actor seeks to maximize expected return while maintaining exploration. For GAC, this objective becomes:
$L_{\text{actor}}(\phi)=\mathbb{E}_{s\sim\mathcal{D}}\left[-\kappa(s)-\min_{i=1,2}Q_{\theta_{i}}(s,\mathbf{a})\right],$ (4)
where $\mathcal{D}$ is the replay buffer, $\phi$ denotes the actor parameters, and $\mathbf{a}$ is generated via GAC’s geometric mechanism in (2) with current state $s$. The term $\kappa(s)$ serves as a learned exploration controller that replaces entropy regularization. Smaller values of $\kappa$ increase the contribution of stochastic noise in the geometric mixing defined in (2), thereby promoting exploration. In contrast, larger values lead to more deterministic actions. This removes the need for temperature tuning in SAC. Unlike traditional entropy-based methods, GAC never computes probability densities. Instead, exploration emerges directly from geometric structure. This design is theoretically justified by directional statistics, where higher concentration naturally corresponds to lower entropy (see Appendix A.4).
The soft Q-function update follows standard SAC with our exploration controller. The target value incorporates the minimum of two Q-networks for stability:
$y(r_{t},s^{\prime})=r_{t}+\gamma\left(\min_{i=1,2}Q_{\theta^{\prime}_{i}}(s^{\prime},\mathbf{a}^{\prime})-\kappa(s^{\prime})\right),$ (5)
where $\theta^{\prime}_{i}$ denotes the parameters of the target Q-network, $s^{\prime}$ is the next state, and $\mathbf{a}^{\prime}$ is generated from $s^{\prime}$ using GAC’s geometric mechanism in (2). Despite replacing Gaussian policies with geometric action generation, GAC maintains the essential properties for convergence in the SAC framework. The bounded action space and smooth geometric operations ensure that the soft Bellman operator remains a contraction (see Appendix A.5 for formal analysis).
3.5 Practical Considerations
Parameter Efficiency. GAC requires only $d+1$ outputs (direction vector plus scalar concentration) compared to $2d$ for diagonal Gaussian policies, where parameter count reduces by nearly 50%.
Computational Efficiency. The sampling procedure involves only normalization and linear interpolation, avoiding rejection sampling or special function evaluations. The computational complexity is $O(d)$ per sample, compared to $O(dk)$ for vMF sampling where $k$ is the expected number of rejections (typically $k\in[2,10]$ for high concentrations).
Hyperparameter Selection. Due to the geometry of high-dimensional spheres, a unit vector in $\mathbb{R}^{d}$ has expected per-dimension magnitude $\mathbb{E}[|\mu_{i}|]\approx 1/\sqrt{d}$, which becomes vanishingly small as $d$ increases (e.g., $\approx 0.24$ for $d=17$). Without scaling, such weak actions would be ineffective for control. The radius $r=2.5$ provides principled rescaling: after spherical mixing with typical $w(\kappa)\approx 0.85$, this yields per-dimension actions around 0.60.60.90.9, which fall comfortably within the normalized action bounds [1,1][-1,1] and ensure effective actuation. While this default works robustly across diverse tasks, certain environments benefit from adjusted scaling (e.g., Ant-v4’s multi-leg coordination performs better with $r=1.0$ for finer control).
4 Experiments
4.1 Experimental Setup
Environments. We evaluate GAC across six standard MuJoCo (Todorov et al., 2012) continuous control benchmarks that collectively test different aspects of control: HalfCheetah-v4 (speed optimization), Ant-v4 (multi-leg coordination), Humanoid-v4 (complex balance), Walker2d-v4 (bipedal locomotion), Hopper-v4 (unstable dynamics), and Pusher-v4 (object manipulation). These environments span action dimensions from 3 to 17 and present diverse challenges from optimization to precise coordination.
Baselines. We compare against SAC (Gaussian + $\tanh$), TD3 (deterministic + noise), and PPO (clipped objectives), using recommended hyperparameters from CleanRL (Huang et al., 2022). All implementations are standardized for fair comparison. See Appendix C.1 for details.
Training Protocol. All algorithms are trained for 1M environment steps with 8 parallel environments, except Pusher-v4, which runs for 500K steps due to faster convergence. We use 5 random seeds {0, 10, 42, 77, 123} and report mean episodic returns $\pm$ standard deviation.
4.2 Main Results
Table 1: Average return over final training steps (1M for all tasks, 500K for Pusher-v4). Bold indicates best performance. GAC achieves competitive performance across diverse control tasks.
Environment GAC (Ours) SAC TD3 PPO
Hopper-v4 1952 $\pm$ 285 2094 $\pm$ 604 2896 $\pm$ 749 2118 $\pm$ 124
Walker2d-v4 5165 $\pm$ 334 5152 $\pm$ 608 4457 $\pm$ 457 2874 $\pm$ 517
Pusher-v4 -32 $\pm$0 -23 $\pm$ 2 -27 $\pm$ 1 -78 $\pm$ 9
HalfCheetah-v4 12750 $\pm$ 758 12540 $\pm$ 517 12208 $\pm$ 799 1608 $\pm$ 793
Ant-v4 5633 $\pm$ 158 4094 $\pm$ 1039 3531 $\pm$ 1263 1969 $\pm$ 778
Humanoid-v4 5823 $\pm$ 121 5717 $\pm$ 123 5819 $\pm$ 278 619 $\pm$ 59
Figure 2: Learning curves on (a) Hopper-v4, (b) Walker2d-v4, (c) Pusher-v4, (d) HalfCheetah-v4, (e) Ant-v4, and (f) Humanoid-v4.
Performance Analysis. Table 1 and Figure 2 present our main experimental results across six MuJoCo benchmarks. GAC demonstrates strong performance, achieving the best results on 4 out of 6 tasks and remaining highly competitive on others. As expected, PPO performs substantially worse across all environments, consistent with its known limitations in exploration and entropy scheduling.
High-dimensional Control Excellence. GAC shows particular strength in complex, high-dimensional tasks. On Ant-v4 (8D action space), GAC achieves 5633 $\pm$ 158, significantly outperforming SAC (4094 $\pm$ 1039) by 37.6% and TD3 (3531 $\pm$ 1263) by 59.5%. The large variance in SAC’s performance ($\pm$ 1039) suggests instability in Gaussian policies for multi-leg coordination, while GAC maintains remarkable consistency ($\pm$ 158). On HalfCheetah-v4 (6D), GAC achieves the highest return of 12750 $\pm$ 758, slightly exceeding SAC and outperforming TD3 by 4.4%.
Learning Efficiency. Beyond final performance, GAC exhibits superior learning dynamics. In Figure 2(d-e), GAC shows faster initial learning on Ant environments, reaching near-optimal performance by 200k steps while SAC and TD3 continue improving until 400k steps. This efficiency stems from GAC’s geometric structure eliminating the need for entropy tuning, as the learned $\kappa$ naturally balances exploration and exploitation without manual temperature scheduling.
Trade-off Between Stability and Expressiveness. GAC demonstrates consistent learning stability across environments, achieving the lowest variance on Pusher-v4 ($32\pm 0$) and competitive performance on Walker2d, HalfCheetah, Ant and Humanoid. This stability advantage is particularly pronounced in high-dimensional spaces, where GAC’s bounded geometric structure prevents the exploration instabilities inherent to squashed Gaussian policies.
However, its performance on Hopper-v4 ($1952\pm 285$) lags behind TD3 ($2896\pm 749$), and similarly underperforms on Pusher-v4. We attribute this to a fundamental trade-off introduced by spherical normalization: while it ensures consistent gradient flow and exploration stability, it imposes a geometric prior that assumes the optimal actions lie close to a unit hypersphere. This assumption aligns well with high-dimensional locomotion tasks, where the optimal actions often reside on a spherical shell due to norm concentration effects, but may be suboptimal for tasks involving asymmetric or contact-rich dynamics. For instance, Hopper’s single-leg locomotion requires strong, unbalanced thrusts that extend beyond the unit sphere, and Pusher’s manipulation demands induce anisotropic action distributions better captured by elliptical or non-spherical structures. These observations suggest future directions for adaptive or task-conditioned geometric priors that modulate between spherical, elliptical, or unconstrained action manifolds based on task characteristics.
Challenging Environments. Humanoid-v4 (17D action space) proves challenging for all methods. During training, GAC obtains 5823 $\pm$ 121, TD3 achieves 5819 $\pm$ 278, while SAC achieves 5717 $\pm$ 123. However, post-training evaluation of GAC achieves 6591 $\pm$ 53, suggesting that the algorithm learns high-quality policies while maintaining exploration during training (detailed evaluation protocol in supplementary materials). This validates that spherical normalization effectively handles complex action spaces and balances exploration-exploitation trade-offs.
4.3 Ablation Studies
Figure 3: Ablation study on HalfCheetah-v4. Default GAC ($r=2.5$, adaptive $\kappa$) performs best.
We conduct comprehensive ablation studies to analyze the contribution of each component in GAC. Table 2 and Figure 3 presents results on HalfCheetah-v4 across 5 random seeds, systematically dissecting the effect of target magnitude $r$ and key architectural components.
Target magnitude $r$ critically affects performance. The choice of $r=2.5$ is theoretically motivated by high-dimensional sphere geometry: a unit vector in $\mathbb{R}^{d}$ has expected per-dimension magnitude $\mathbb{E}[|\mu_{i}|]\approx 1/\sqrt{d}$, which becomes vanishingly small as $d$ increases (e.g., $\approx 0.24$ for $d=17$), rendering unscaled actions ineffective for control. Figure 3 empirically validates this insight: reducing $r$ from 2.5 to 1.5 causes catastrophic 43% performance degradation due to limited exploration and actuation. In contrast, increasing $r$ to 3.5 has minimal impact ($4.1\%$), suggesting saturation. The default $r=2.5$, after spherical mixing with typical $w(\kappa)\approx 0.85$, yields per-dimension actions in the range 0.6–0.9. This range falls well within the normalized bounds [1,1][-1,1] and strikes an effective balance for HalfCheetah-v4’s locomotion dynamics.
Table 2: Ablation study of GAC components on HalfCheetah-v4. Results averaged over 5 seeds.
Configuration Final Return Relative Key Observation
GAC (default with $\kappa$ and $r=2.5$) 12750 $\pm$ 758 baseline Optimal balance
Target magnitude ablation:
   $r=3.5$ 12229 $\pm$ 422 -4.1% Slightly over-scaled actions
   $r=1.5$ $72$ $\pm$ 1235 -43.0% Severely limited action range
Component ablation:
   w/o $\kappa$ controller 11370 $\pm$ 643 -10.8% No adaptive exploration
   w/o normalization Diverged N/A Gradient explosion at 5k steps
   Raw action output Collapsed N/A Unbounded actions, NaN loss
Adaptive exploration via $\kappa$ is essential. Removing the learnable $\kappa$ controller while maintaining $r=2.5$ reduces performance by 10.8% and slows convergence. The learned $\kappa$ enables state-dependent exploration by providing high concentration in confident states while maintaining diversity in uncertain regions. This adaptive behavior emerges naturally from the value-based objective without explicit curriculum or scheduling. Notably, this mechanism provides an elegant alternative to entropy-based regularization: rather than maximizing entropy uniformly, GAC modulates exploration geometrically through directional mixing. This not only simplifies the optimization pipeline by eliminating entropy terms, but also improves interpretability, as $\kappa(s)$ can be viewed as a soft confidence score indicating how deterministic the policy should be at a given state. As a result, exploration becomes structure-aware and implicitly guided by the task dynamics.
Geometric structure ensures stability. As shown in Table 2, removing normalization leads to divergence within 5k steps due to unbounded gradients, as output norms grow without spherical projection and trigger explosion. Likewise, using raw unbounded outputs collapses immediately (NaN losses) as actions exceed environment limits and destabilize training. These ablations confirm that GAC’s spherical geometry is essential for stability. Beyond preventing numerical issues, the constraint shapes a consistent optimization landscape where all actions share equal norm, removing scale ambiguity and acting as an implicit regularizer against degenerate solutions.
5 Conclusion
This work demonstrates that effective continuous control does not require complex probability distributions. GAC achieves competitive or superior performance across diverse benchmarks using a single geometric operation: $\mathbf{a}=r\cdot\text{normalize}(w(\kappa)\cdot\bm{\mu}+(1-w(\kappa))\cdot\bm{\xi})$. By replacing distributional modeling with direct geometric mixing, we reduce parameter count by nearly 50% and improve performance, achieving a 37.6% improvement over SAC on Ant-v4 and the best results on 4 out of 6 MuJoCo benchmark tasks.
The success of GAC validates a broader principle: respecting the geometric structure of action spaces can be more effective than sophisticated probabilistic machinery. The consistent performance with a fixed radius $r=2.5$ across diverse tasks suggests that learning correct action directions is often the primary challenge, while magnitude can be addressed with simple task-specific constants.
Our method eliminates the computational burden of density calculations, Bessel functions, and rejection sampling, while avoiding the gradient pathologies of $\tanh$-squashed Gaussians. The learnable concentration parameter $\kappa$ provides adaptive exploration without explicit entropy computation, demonstrating that exploration-exploitation balance can emerge from geometric structure rather than information-theoretic regularization.
Limitations and Future Work. While GAC demonstrates strong empirical performance, several avenues remain for investigation. The action radius $r$, though robust across tasks (with $r=2.5$ working well for most benchmarks), could potentially be learned adaptively for task-specific optimization. The theoretical connection between our geometric exploration mechanism and information-theoretic quantities, while empirically validated through strong performance, warrants deeper mathematical analysis. Additionally, extending the geometric approach to discrete or hybrid action spaces presents exciting challenges for general-purpose control.
Despite these open questions, GAC’s success suggests that the Geometric Simplicity Principle—replacing probabilistic complexity with geometric structure—could transform other areas of RL. Future work might explore geometric approaches to value function approximation, hierarchical control, or multi-agent coordination. By showing that simple geometric operations can replace complex probabilistic frameworks, GAC challenges the prevailing assumption that sophisticated distributions are necessary for continuous control and opens new directions in geometric RL. Ultimately, control is not about predicting densities, but about choosing directions. GAC shows that when geometry is respected, simplicity is not a compromise—but a strength. We hope this work inspires further efforts to rethink RL through the lens of geometric structure, not just statistical modeling.
Appendix A Theoretical Analysis
A.1 Proof of Theorem 1
Proof.
Consider the unnormalized mixture vector:
$\mathbf{v}=w\bm{\mu}+(1-w)\bm{\xi},$ (6)
where $\bm{\mu}\in\mathbb{S}^{d-1}$ is the deterministic mean direction, $\bm{\xi}\sim\text{Uniform}(\mathbb{S}^{d-1})$ is uniform spherical noise, and $w=w(\kappa)\in[0,1]$ is the mixing weight.
Due to the symmetry of the uniform distribution on $\mathbb{S}^{d-1}$, any uniform random vector on the sphere has zero expectation:
$\mathbb{E}_{\bm{\xi}}[\bm{\xi}]=\mathbf{0}.$ (7)
Therefore, the expectation of the mixture vector is:
$\mathbb{E}_{\bm{\xi}}[\mathbf{v}]=\mathbb{E}_{\bm{\xi}}[w\bm{\mu}+(1-w)\bm{\xi}]=w\bm{\mu}+(1-w)\mathbb{E}_{\bm{\xi}}[\bm{\xi}]=w\bm{\mu}.$ (8)
This result holds exactly for any dimension $d\geq 2$. ∎
A.2 Concentration Analysis
While Theorem 1 characterizes the expected direction of the unnormalized mixture, it does not capture how tightly the samples are concentrated around this direction. We therefore analyze the cosine similarity between normalized samples and the mean direction to quantify concentration. Let $\hat{\mathbf{v}}=\mathbf{v}/\|\mathbf{v}\|_{2}$ denote the normalized mixture, where $\|\cdot\|_{2}$ is the L2 norm.
The cosine similarity between $\hat{\mathbf{v}}$ and $\bm{\mu}$ measures directional alignment:
$\cos\angle(\hat{\mathbf{v}},\bm{\mu})=\frac{\mathbf{v}^{\top}\bm{\mu}}{\|\mathbf{v}\|_{2}}.$ (9)
For the numerator:
$\mathbf{v}^{\top}\bm{\mu}$ $=(w\bm{\mu}+(1-w)\bm{\xi})^{\top}\bm{\mu}$ (10)
$=w\bm{\mu}^{\top}\bm{\mu}+(1-w)\bm{\xi}^{\top}\bm{\mu}$
$=w\|\bm{\mu}\|^{2}+(1-w)\bm{\xi}^{\top}\bm{\mu}$
$=w+(1-w)\bm{\xi}^{\top}\bm{\mu},$
where the last step uses $\|\bm{\mu}\|^{2}=1$ since $\bm{\mu}\in\mathbb{S}^{d-1}$.
Taking expectations over the uniform distribution:
$\mathbb{E}_{\bm{\xi}}[\mathbf{v}^{\top}\bm{\mu}]=w+(1-w)\mathbb{E}_{\bm{\xi}}[\bm{\xi}^{\top}\bm{\mu}]=w,$ (11)
since $\mathbb{E}_{\bm{\xi}}[\bm{\xi}^{\top}\bm{\mu}]=0$, as a fixed unit vector and a random unit vector on the sphere are uncorrelated in expectation under uniform sampling.
To estimate the norm, we assume that the inner product $\bm{\mu}^{\top}\bm{\xi}$ remains small, which is typically the case when $\bm{\xi}$ is uniformly sampled from the sphere. This leads to the approximation:
$\|\mathbf{v}\|_{2}^{2}=\|w\bm{\mu}+(1-w)\bm{\xi}\|^{2}=w^{2}+(1-w)^{2}+2w(1-w)(\bm{\mu}^{\top}\bm{\xi})\approx w^{2}+(1-w)^{2}.$ (12)
This simplification holds well in practice as confirmed by our empirical results (see Appendix A.3).
For action spaces ($d\gg 1$), concentration of measure implies that $\bm{\xi}^{\top}\bm{\mu}$ concentrates tightly around zero with high probability. Under this regime:
$\mathbb{E}[\cos\angle(\hat{\mathbf{v}},\bm{\mu})]\approx\frac{w}{\sqrt{w^{2}+(1-w)^{2}}}.$ (13)
For typical operating ranges where $w\in[0.6,0.99]$ (corresponding to $\kappa\in[0.5,5]$), this quantity closely tracks $w$ itself. For example, when $w=0.9$, the ratio equals 0.994, validating our use of $w(\kappa)$ as an effective concentration parameter.
A.3 Empirical Validation
Table A.1: Concentration control validation: theoretical vs. measured. The close match between measured concentration and theoretical $w$ validates Theorem 1.
$\kappa$ Weight $w$ Concentration Angle Std. ()
-2.0 0.119 0.091 39.3
-1.0 0.269 0.253 36.3
0.0 0.500 0.678 19.7
0.5 0.622 0.875 9.1
1.0 0.731 0.956 5.0
2.0 0.881 0.994 1.7
Figure A.1: 3D visualization of GAC sample distributions for $\kappa\in\{-2,0,0.5,1\}$. Arrows indicate target direction $\bm{\mu}$. Colors represent cosine similarity with $\bm{\mu}$ (blue=low, red=high). Higher $\kappa$ values produce more concentrated distributions.
To complement our theoretical analysis, we empirically verify that the concentration parameter $\kappa$ provides direct and monotonic control over the sample distribution’s properties. For each $\kappa\in\{-2,-1,0.0,0.5,1.0,2.0\}$, we generate 500 samples using the GAC mechanism and measure their key concentration metrics, summarized in Table A.1 and visualized in Figure A.1.
Metrics Explanation:
  • Weight $w(\kappa)$: The mixing weight $w(\kappa)=\sigma(\kappa)$ that controls interpolation between deterministic direction and uniform noise. This is deterministic given $\kappa$.
  • Concentration: The empirical mean cosine similarity between normalized samples and the target direction $\bm{\mu}$, measuring how aligned the final actions are with the intended direction.
[…]
Empirical Validation. We empirically validate the theoretical result in Theorem 1 by measuring the directional concentration of sampled actions under different $\kappa$ values. Results show strong agreement between measured and theoretical concentration (correlation $\approx$ 0.95), with angular standard deviation decreasing monotonically from 39.3 at $\kappa=-2$ (high exploration) to 1.7 at $\kappa=2$ (near-deterministic). These results confirm that GAC achieves precise, vMF-like concentration control through simple geometric operations—without requiring modified Bessel functions or rejection sampling. The monotonic relationship between $\kappa$ and directional concentration further validates our approach to structured exploration.
A.4 Exploration Control and Entropy Connection
A.4.1 Exploration Control Mechanism
GAC’s concentration parameter $\kappa$ functions as an adaptive exploration controller that learns when to explore versus exploit based on the value landscape. Unlike traditional entropy regularization that requires computing probability densities, $\kappa$ directly modulates the geometric mixing between deterministic and stochastic components. While GAC implicitly defines a mixture distribution through this geometric operation, it crucially avoids any density or entropy computation during optimization. The effectiveness of $\kappa$ as an exploration signal emerges from three complementary perspectives:
1) Geometric Perspective. As $\kappa$ increases, the mixing weight $w(\kappa)=\sigma(\kappa)$ increases monotonically from 0 to 1, causing samples to concentrate progressively around $\bm{\mu}$. Empirically, the angular standard deviation decreases from 39.3 at $\kappa=-2$ to 1.7 at $\kappa=2$ (Table A.1), directly demonstrating decreasing distributional uncertainty.
2) Information-Theoretic Perspective. For spherical distributions, concentration and entropy are fundamentally inversely related. The vMF distribution provides a theoretical benchmark:
$\mathcal{H}_{\text{vMF}}=\log C_{d}(\kappa)-\kappa A_{d}(\kappa)\approx-\kappa+\frac{d-1}{2}\log\kappa+\text{const},$ (14)
where $C_{d}(\kappa)$ is the normalization constant and $A_{d}(\kappa)$ the mean resultant length; the approximation holds for large $\kappa$, with the dominant linear term $\kappa$ justifying our exploration term as higher-order terms contribute little in practice.
Note: The relationship $\mathcal{H}\approx-\kappa$ serves as conceptual motivation rather than rigorous derivation. GAC uses $\kappa$ directly as an exploration signal without ever computing actual entropy. This approximation provides theoretical intuition for why $\kappa$ effectively balances exploration-exploitation, but the method’s success does not depend on this mathematical correspondence.
3) Empirical Validation. Our experiments confirm that $\kappa$ effectively captures exploration pressure:
  • The correlation between $w(\kappa)$ and measured concentration exceeds 0.95 (Figure A.1)
  • As $\kappa$ increases: $w(\kappa)=\sigma(\kappa)\to 1$ (more deterministic/exploitative)
  • As $\kappa$ decreases: $w(\kappa)=\sigma(\kappa)\to 0$ (more stochastic/exploratory)
  • GAC with $\kappa$ achieves superior performance across all benchmarks (Table 1)
  • $\kappa$’s smooth effect on exploration supports stable policy optimization
A.4.2 Theoretical Connection to Entropy
While GAC operates without computing distributions, the exploration controller $\kappa$ exhibits a natural connection to entropy in directional statistics. For vMF distributions on the unit sphere, the differential entropy is given by:
$\mathcal{H}_{\text{vMF}}=\log C_{d}(\kappa)-\kappa A_{d}(\kappa)\approx-\kappa+\frac{d-1}{2}\log\kappa+\text{const},$ (15)
where $C_{d}(\kappa)=\frac{\kappa^{d/2-1}}{(2\pi)^{d/2}I_{d/2-1}(\kappa)}$ is the normalization constant and $A_{d}(\kappa)=\frac{I_{d/2}(\kappa)}{I_{d/2-1}(\kappa)}$ is the mean resultant length, with $I_{v}$ denoting the modified Bessel function of the first kind. The asymptotic behavior of $A_{d}(\kappa)$ is:
$A_{d}(\kappa)\approx 1-\frac{d-1}{2\kappa}+\mathcal{O}(\kappa^{-2}).$ (16)
Substituting into $\mathcal{H}_{\text{vMF}}$ and simplifying, we obtain:
$\mathcal{H}_{\text{vMF}}\approx-\kappa+\frac{d-1}{2}\log\kappa+\text{const}.$ (17)
The leading term $\kappa$ dominates, with logarithmic corrections diminishing in relative magnitude for practical $\kappa$ values. This validates using $\kappa$ as an exploration controller that faithfully reflects the tradeoff between concentration and uncertainty. While GAC does not explicitly follow the vMF distribution, it inherits the same qualitative dependency between $\kappa$ and sample concentration through its spherical interpolation mechanism. This connection validates why $\kappa$ effectively balances exploration-exploitation in the SAC framework. Empirical results in Appendix A.3 further confirm the effectiveness of this surrogate.
Key distinction: While this mathematical relationship exists, GAC fundamentally differs from entropy-regularized methods:
  • Traditional SAC: Computes $\mathcal{H}[\pi]=-\mathbb{E}[\log\pi(\mathbf{a}|s)]$ requiring explicit densities
  • GAC: Uses $\kappa$ as exploration signal without any distributional computation
This allows GAC to achieve entropy-like regularization benefits through purely geometric operations, eliminating computational overhead while maintaining theoretical grounding.
A.5 SAC Convergence with GAC
Theorem 2: GAC maintains key properties required for SAC-style convergence under standard regularity conditions.
Remark: We provide a sketch of the key arguments. A rigorous convergence proof would require extensive measure-theoretic analysis beyond the scope of this work. Our empirical results across diverse environments provide strong evidence for convergence in practice.
Proof Sketch.
We establish that GAC maintains the key properties required for SAC convergence.
Soft Bellman Contraction. The soft Bellman operator with GAC takes the form:
$\mathcal{T}^{\pi}Q(s,\mathbf{a})=r_{t}(s,\mathbf{a})+\gamma\mathbb{E}_{s^{\prime}\sim p(\cdot|s,\mathbf{a})}[V^{\pi}(s^{\prime})],$ (18)
where the soft value function incorporates our exploration controller:
[…]
The contraction property requires:
(i) Continuity: GAC’s action generation is continuous in parameters:
  • Direction mapping $\bm{\mu}(s)=f_{\mu}(s)/\|f_{\mu}(s)\|_{2}$ is continuous for $f_{\mu}(s)\neq 0$
  • Mixing weight $w(\kappa)=\sigma(\kappa)$ is $C^{\infty}$ smooth
  • Action normalization ensures $\|\mathbf{a}\|=r$ (bounded)
(ii) Regularization: The exploration controller provides consistent exploration pressure:
$\nabla_{\kappa}[-\kappa]<0,$ (20)
encouraging exploration when concentration becomes excessive.
Under these conditions and assuming bounded rewards $|r(s,\mathbf{a})|\leq R_{\max}$, the operator $\mathcal{T}^{\pi}$ is a $\gamma$-contraction:
$\|\mathcal{T}^{\pi}Q_{1}-\mathcal{T}^{\pi}Q_{2}\|_{\infty}\leq\gamma\|Q_{1}-Q_{2}\|_{\infty}.$ (21)
Policy Improvement. GAC approximates policy improvement through gradient-based optimization:
$L_{\text{actor}}=\mathbb{E}_{s\sim\mathcal{D},\bm{\xi}}[-\kappa(s)-Q(s,\mathbf{a})],$ (22)
where $\mathbf{a}$ is generated using GAC’s mechanism with random noise $\bm{\xi}$. This objective drives the policy toward high-value regions (via the Q-term) while maintaining exploration (via the $\kappa$ term), achieving similar goals to SAC’s entropy-regularized policy improvement without explicit distributional computations.
Convergence Benefits. GAC improves convergence through three key design choices:
  • Stable gradients: Spherical normalization avoids $\tanh$-induced saturation, preserving direction gradients.
  • Bounded actions: All actions satisfy $\|\mathbf{a}\|=r$, preventing value divergence from out-of-bound actions.
  • Adaptive exploration: The learnable $\kappa$ balances exploration and exploitation without external schedules.
Exploration–Exploitation Balance. The loss function induces an implicit tradeoff between exploration and exploitation. The direct gradient $\partial\mathcal{L}/\partial\kappa=-1$ promotes exploration by reducing $\kappa$, increasing action noise. In contrast, the Q-value term encourages exploitation: when higher $\kappa$ leads to better actions (and thus higher $Q$), gradients through $Q(s,\mathbf{a}(s;\kappa))$ push $\kappa$ upward. This dynamic balance emerges naturally from actor–critic interplay, without explicit entropy terms or temperature tuning.
Natural Stabilization of $\kappa$. Despite being unconstrained, $\kappa$ remains bounded ($\in[0,5]$ empirically) through:
  • Sigmoid saturation: $w(\kappa)=\sigma(\kappa)$ flattens for large $\kappa$, capping its effect.
  • Gradient feedback: High $\kappa$ leads to deterministic actions; in noisy environments, this reduces $Q$, discouraging overconfidence.
  • No explicit clipping: Yet $\kappa$ stabilizes naturally via loss dynamics.
These mechanisms ensure convergence stability without manual regularization.
Appendix B Gradient Flow Analysis in Tanh-Squashed Policies
Figure A.2: Distribution of pre-squashed Gaussian samples from a trained SAC policy. Red areas indicate saturated gradients ($|\tanh^{\prime}(x)|<0.05$), with 46.4% of samples falling into these regions. Dashed lines show the [1,1][-1,1] $\tanh$ boundaries. This mismatch between unbounded Gaussians and bounded action spaces motivates GAC’s direct geometric approach.
[…]
We sampled pre-squashed actions from the SAC policy throughout training, particularly during the stable performance phase near convergence (0.8M–1.0M steps).Figure A.2 visualizes the distribution of these raw actions before squashing, color-coded by gradient magnitude.
Key Observations:
  • A substantial fraction (46.4%) of pre-squashed samples have gradients below 0.1, indicating severe saturation.
  • The distribution exhibits heavy tails beyond $|\tilde{\mathbf{a}}|>2.5$, where gradient flow is minimal.
  • SAC compensates through entropy regularization ($\alpha\approx 0.2$), which maintains exploration diversity despite reduced gradients.
This analysis does not imply SAC is ineffective, as it remains highly successful in practice. Rather, it highlights a structural inefficiency in that significant computational effort is spent managing the mismatch between unbounded distributions and bounded spaces. GAC sidesteps this issue entirely by operating directly on the unit sphere, ensuring consistent gradient flow without requiring squashing or entropy-driven exploration.
Note: This visualization shows pre-squashed samples. The actual gradient flow during training is modulated by reparameterization and entropy regularization, which prevent complete saturation. However, this analysis does not diminish SAC’s effectiveness but highlights an opportunity for geometric alternatives like GAC that avoid this structural inefficiency entirely.
Appendix C Reproducibility
To ensure complete reproducibility and facilitate future research, we provide comprehensive implementation details and open-source resources for immediate verification of our results.
C.1 Implementation Details
Network Architecture: GAC uses a shared backbone with separate heads for direction and concentration:
  • Backbone: Linear(obs_dim, 256) → ReLU → Linear(256, 256) → ReLU
  • Direction head: Linear(256, action_dim)
  • Concentration head: Linear(256, 64) → ReLU → Linear(64, 1)
Key Hyperparameters:
  • Learning rates: $3\times 10^{-4}$ (actor), $1\times 10^{-3}$ (critic)
  • Batch size: $256$, Buffer size: $10^{6}$
  • Target network update: $\tau=0.005$
  • Discount factor: $\gamma=0.99$
  • Action radius $r$: $2.5$ for most tasks, $1.0$ for Ant-v4
We adopt standard hyperparameters from CleanRL (Huang et al., 2022) without task-specific tuning, highlighting that GAC’s gains stem from structural design rather than careful optimization.
C.2 Computational Efficiency
Computational Efficiency: GAC’s sampling requires only: 1). Two forward passes (direction and concentration networks); 2). One normalization operation; 3). One linear interpolation; 4). One final scaling. This yields approximately 6× speedup compared to vMF sampling with rejection methods.
C.3 Code and Model Release
To enable immediate verification of our results, we provide the following:
Pre-trained Models: Trained policy weights demonstrating reported performance.
Evaluation Scripts: Evaluation protocols enabling exact reproduction of reported metrics.
Availability: Anonymized resources available during review at:
https://github.com/geometric-rl-anonymous/Geometric-Action-Control-for-Reproducibility
, including the trained models and evaluation scripts for the primary benchmarks (Walker2d-v4, HalfCheetah-v4, Ant-v4, and Humanoid-v4) used in our analyses, to facilitate easy verification and reproduction of our results by reviewers. A complete open-source release is planned upon paper acceptance with comprehensive documentation and examples.